OpenStudy (anonymous):

Using an alternate method, explain how to compute: a) 1/2 + 1/3 b) 1/2 x 1/3

OpenStudy (isaiah.feynman):

\[\frac{ a }{ b }+\frac{ c }{ d } = \frac{ ad+bc }{ bd }\] \[\frac{ a }{ b } \times \frac{ c }{ d } = \frac{ ac }{ bd }\]